### v2

# AttributionScanner: A Visual Analytics System for Model Validation with Metadata-Free Slice Finding

Abstract
Data slice finding is an emerging technique for validating machine learning (ML) models by identifying and analyzing subgroups in a dataset that exhibit poor performance, often characterized by distinct feature sets or descriptive metadata. However, in the context of validating vision models involving unstructured image data, this approach faces significant challenges, including the laborious and costly requirement for additional metadata and the complex task of interpreting the root causes of underperformance. To address these challenges, we introduce AttributionScanner, an innovative human-in-the-loop Visual Analytics (VA) system, designed for metadata-free data slice finding. Our system identifies interpretable data slices that involve common model behaviors and visualizes these patterns through an Attribution Mosaic design. Our interactive interface provides straightforward guidance for users to detect, interpret, and annotate predominant model issues, such as spurious correlations (model biases) and mislabeled data, with minimal effort. Additionally, it employs a cutting-edge model regularization technique to mitigate the detected issues and enhance the model’s performance. The efficacy of AttributionScanner is demonstrated through use cases involving two benchmark datasets, with qualitative and quantitative evaluations showcasing its substantial effectiveness in vision model validation, ultimately leading to more reliable and accurate models.
Index Terms:
Model validation, visual analytics, data slicing, data-centric AI, human-assisted AI.
I Introduction
As machine learning (ML) technologies continue to prevail in various domains, the importance of model interpretation and validation has grown significantly. Identifying a model’s failure and understanding the underlying reasons are crucial in promoting greater model accuracy, transparency, and accountability. A powerful technique supporting this effort is data slice finding [1], aiming to identify problematic data subgroups, referred to as slices, which often arise from under-represented samples or biased datasets [2]. These data slices are commonly defined by descriptive metadata, with contextual details of the model’s input such as feature descriptions, data collection details, and other relevant information.
After problematic data slices are identified, the remedial strategies often involve model training or fine-tuning with these slices to bolster the model’s overall reliability [3]. To enhance the trustworthiness, eXplainable Artificial Intelligence (XAI) [4] are also utilized in recent techniques to reveal potential causal factors adversely impacting the model and apply targeted model regularization [5] to improve the performance.
Despite the promising benefits of these techniques, several hurdles remain unaddressed for a thorough model validation.
Challenges of Data Slice Finding. State-of-the-art methods for data slice finding demand extensive metadata associated with each instance. Many works use pre-existing metadata collected together with the dataset [6, 7, 8], and others apply pre-trained vision-language models to generate descriptive data [3]. However, there are challenges hindering their applications—On the one hand, it requires significant effort to collect descriptive metadata. On the other hand, trained with general-purpose data, vision-language models can generate erroneous results in domain-specific scenarios. In situations where neither metadata nor suitable pre-trained models are available, obtaining interpretable data slices becomes challenging: requires manual examination of numerous instances or computationally expensive model training [9, 10].
Challenges of XAI for Vision Models. If troublesome data slices have been identified, it also requires non-trivial efforts to uncover the root causes of a mode’s failure. Even with the advanced human-friendly XAI techniques, there still exist a few challenges. Predominantly, most model explanation methods are instance-based, conducting a thorough examination of all instances to derive conclusions based on a big picture is practically unfeasible [11, 12, 13, 14]. Moreover, the insights we can obtain from existing methods are often not actionable, which means what the common failure patterns are and how to address them remain vague and demands further study.
Our approach. We present AttributionScanner, a Visual Analytics (VA) system designed to address the above challenges, providing an interpretable vision model validation framework with metadata-free slice finding. Our innovative workflow employs attribution-based explanation techniques to create interpretable feature representations for slice finding, sorting through unstructured image data into slices with common model behaviors. Moreover, we explain the discriminating features of each data slice via visual summaries, called Attribution Mosaic, to guide users in explaining model’s behaviors and detecting potential issues, thus facilitating model validation. Additionally, AttributionScanner integrates a human-in-the-loop approach to rank and annotate the problematic data slices based on common error types such as mislabeled samples and spurious correlations. Our approach enables users to apply these annotated slices to enhance the model, by correcting the identified errors on either the model or data side. Our main contributions include:
  1. AttributionScanner, a novel framework and VA system to validate vision models with data slice finding. Unlike existing methods, AttributionScanner removes the resource-intensive need for metadata to generate interpretable data slices.
  2. A design for slice summarization, Attribution Mosaic, illustrating the significant pattern of model behaviors in data slices, which reduces the effort in conventional approaches of examining a large volume of individual data samples.
  3. An iterative workflow supporting model validation and enhancement, which ensures generated insights are actionable, enabling users to swiftly detect, understand, and fix model issues related to spurious correlations and mislabeled data.
  4. Qualitative and quantitative evaluation. We underscore the efficacy of AttributionScanner through two use cases with qualitative and quantitative evaluation, highlighting its ability to identify and mitigate hidden model issues.
This paper is structured as follows: in Sec. II, we review related work on data slice finding and XAI that is relevant to our work Sec. III outlines the methodology and system design of AttributionScanner. In Sec. IV, we illustrate two use cases showcasing how AttributionScanner can be utilized for vision model validation. An evaluation of our system and experts’ feedback are discussed in Sec. V. Lastly, we present our future work and conclusions in Sec. VII.
II Related Work
Our method, AttributionScanner, leverages insights from both Data Slice Finding and XAI techniques to pinpoint problematic data slices in vision-based models. This provides a brief overview of the relevant literature on these two topics.
II-A Data Slice Finding
Data slice finding technique is used to identify data subsets where ML models underperform. These subsets are usually defined by a combination of feature-value pairs extracted from tabular metadata. However, the exponential search space associated with this task makes it a challenging problem. To overcome this challenge, heuristics have been presented to approximate the solution. Barash et al. [15] introduced a Combinatorial Modeling-based approach that enumerates the space of value combinations in tabular data up to a set number of parameters. Similarly, Slice Finder [6] uses a decision tree heuristic to identify interpretable data slices that can be described by tree rules. Despite the advantages of heuristics, they may miss important data slices. Therefore, new techniques have been developed to enable the exhaustive search of data slices using novel formulations and parallel architectures. DivExplorer [1] employs frequent pattern mining to enumerate all candidate slices in the data and identify the most relevant ones based on frequency or metric divergence. In contrast, SliceLine [7] introduces a vectorized formulation for slice finding, which can be easily deployed in parallel devices.
The aforementioned methods necessitate structured metadata to segment datasets into meaningful subgroups, a requirement challenging to meet, especially in domains with large volumes of image or text data where detailed meta-information may be scarce. To address this, a prevalent approach involves encoding data samples into the model’s latent space, followed by clustering to group the data and pinpoint problematic slices. For instance, GEORGE [9] employs over-clustering alongside Group Distributionally Robust Optimization [16] to automatically detect and rectify distribution shift-induced errors. Spotlight [10] adopts a soft-clustering strategy, evaluating performance metrics per cluster and pinpointing clusters with the highest performance drops as data slices for user review. Similarly, UDIS [17] implements hierarchical clustering and further GradCAM [12] visual explanations for data slice examination. VisCUIT [18] allows a deeper exploration of the UDIS data slices by enabling slice comparison and the exploration of neuron activations per slice. Although these methods can support the identification of problematic slices, users have to delve into individual samples to understand model failure reasons. Domino [3] mitigates this by utilizing a pre-trained vision-language model for generating textual descriptions of slices, encoding images into model embedding, and clustering them based on shared visual attributes. However, this method’s reliance on pre-trained models for slice identification and explanation can backfire if there’s a significant domain mismatch between the dataset being evaluated and the pre-trained model, possibly yielding unmeaningful slices. In Sec. III, we introduce a methodology capable of producing explainable slices independently of such pre-trained models.
Visualization has also proven effective in identifying problematic subgroups of data. Kahng et al. introduced a data cube analysis method [19], in which the performance of models was assessed on data slices defined by all pairwise combinations of data features. Meanwhile, FairVis [20] enables users to manually slice and inspect data slices based on domain knowledge to identify potential model biases. Similarly, Errudite [21] came up with a domain-specific language to allow the data slicing of textual documents based on linguistic features and model testing under specific conditions. Although manually slicing the data is useful for model analysis, this approach is not scalable since there could be an exponential number of slices to evaluate. To address this issue, SliceTeller [2] combined an automatic slice finding algorithm with a visual analytics tool to allow the analysis and iterative refinement of models. However, it still required structured metadata in order to produce meaningful data slices. In Sec. III, we present a visual analytics tool that requires no metadata or prior information to produce meaningful data slices.
Figure 1: AttributionScanner workflow involves three phases: Explainable Data Slice Finding, Slice Summarization and Annotation, and Slice Error Mitigation. The first phase: GradCAM is used to assist the generation of feature vectors and then obtain data slices. The second phase: users can identify and annotate slice error types such as core/spurious correlations and noisy labels with the help of Attribution Mosaic and Spuriousness propagation. The third phase: the annotation and user-verified Spuriousness are used on the ML side to mitigate slice errors, including (1) Data re-labeling if considered necessary, and (2) noise corruption and Core Risk Minimization (CoRM) applied if spurious correlations are detected.
II-B XAI Techniques for Neural Networks
Given the widespread adoption of deep learning models, there is significant interest in understanding their prediction processes. In particular, visualization-based XAI methods have gained popularity due to their ability to produce intuitive and meaningful results and be applied to explain various deep Convolutional Neural Networks (CNNs). One genre is localization-based explanations, namely attribution heatmaps, which highlight image regions mostly contributing to a specified class prediction, providing class-discriminative localization to support model interpretation. Class Activation Maps (CAM) [11] is a pioneering method, but it requires a specific network architecture, namely the global average pooling layer, to generate visualizations. GradCAM [12] addresses this limitation by using gradients to compute model attributions. In our work, we utilize GradCAM attribution heatmaps in two ways: firstly, we use them to explain model predictions on single images. Secondly, we interpolate the attributions into weight matrices at the latent space to obtain attribution-weighted feature vectors, which are used to obtain and visualize data slices with similar model attribution patterns.
Another popular approach for understanding model behavior is optimization-based visual explanations. Feature Visualization [22, 13] creates images that maximally activate specific neurons, channels, or layers of a neural network, which focuses on understanding features that a model is sensitive to. Feature inversion [23] seeks to understand the information encapsulated in a group of spatial feature vectors. Specifically, it visually elucidates the shared patterns significant in those latent vectors from the model’s perspective.
III AttributionScanner
In this section, we introduce AttributionScanner, a novel approach for interpretable ML model validation with metadata-free, data slice-level analysis. To foster a thorough understanding of our system, we first outline the requirements that shaped our design choices. Then, we delve into the visual components of AttributionScanner that enable interpretable model validation. We then detail our slice computation and summarization method, which allow us to generate data slices with consistent model attribution without additional metadata. Finally, we show how we leverage the insights obtained from our system to enhance the model.
III-A Design Requirements
Slice finding has garnered increased attention due to its potential to facilitate a thorough ML model evaluation. However, this task is particularly challenging in the realm of vision models considering the inherent unstructured nature of vision data. A pivotal shortcoming we observed in existing approaches is their dependency on metadata, including pre-collected or vision-language model-generated ones, to compute meaningful data slices. However, the acquisition of metadata is resource-intensive, demanding substantial human effort and financial investment. Additionally, the vision-language models are commonly trained on general-purpose datasets, potentially leading to inaccurate or biased results in specific domains. Stemming from these observations, we delineated the subsequent requirements for a system capable of conducting metadata-free, slice-driven model validation:
  1. R1.
    Metadata-free. Our method should yield meaningful data slices without metadata, i.e., descriptive information other than the model’s training data.
  2. R2.
    Interpretability. We should provide sufficient and intuitive information to enhance model transparency. Diverging from prior metadata-free approaches that lack explanation capability, our system should enable both intuitive explanations of a data slice’s shared attributes and inspection of individual instances for insight verification as needed.
  3. R3.
    Slice Overview. Mandatory manual inspection of individual images is impractical for model validation, especially with a high volume of data. To enhance efficiency, our system should provide a visual summary of data slices, depicting each data slice’s key features and attributions.
  4. R4.
    Actionable insights. The system should ensure the obtained insights are actionable, which requires careful design of visual guidance, issue annotation, and corresponding model regularization. The framework should allow for further expert engagement, such as re-labeling or model re-training, to enhance model performance.
III-B System Workflow
AttributionScanner is a human-in-the-loop system, taking as input an image dataset and a trained CNN for image classification, and yielding interpretable data slices to assist human experts in metadata-free vision model validation.
A workflow of our approach is presented in Fig. 1, where we foster explainable slice finding without metadata ( R1, R2) by leveraging XAI technique GradCAM [12] and Feature Inversion [13]. The first “explainable data slice finding” phase interpolates model attributions from GradCAM into the latent space to craft attribution-weighted feature vectors, which enable us to form an attribution representation space with locally consistent model attributes ( R1). Then we conduct clustering over this space to generate data slices. Although the produced slices maintain consistent attributions, they still demand significant human effort to identify and understand their shared patterns. To streamline slice exploration and issue detection, the “slice summarization and annotation” phase introduces Attribution Mosaic, which visually elucidates each slice’s dominant attributions ( R3), synchronized with instance-level heatmaps for insight verification ( R2). Upon user annotation of an issue like “spurious correlations”, our Spuriousness Propagation method automatically estimates a spuriousness score for each slice to help users uncover other hidden model biases ( R4). Afterward, our workflow enables the mitigation of detected issues in data or model, corresponding to the “slice error mitigation” phase ( R4).
The rest of this section delineates our interface for data slicing-based model validation, explains our presented data slicing and summarization approach, discusses design considerations, and demonstrates how the model can be fine-tuned to rectify detected slice issues.
Figure 2: AttributionScanner applied to the model validation of a hair color classifier trained on the CelebA dataset. A System Menu, enabling the selection of dataset, model, and visualization options (Layout, Confusion Matrix, Scatter Plot/Attribution Mosaic, and Contour Visibility). B Slice Table, showing slice metrics. C Attribution Mosaic, showing a visual overview of all data slices, which can also be displayed as a confusion matrix view ( E ). D Slice Detail View, showing individual images or Attribution Heatmaps belonging to a selected data slice.
III-C System Interface
AttributionScanner is comprised of four main components. Fig. 2 depicts a use case where our system is used to validate a hair color classification model.
The System Menu (Fig. 2  A) provides options for dataset and model selection, visualization layout configuration, and color encoding choice. Two visualization layouts are available: Combination view and Confusion Matrix view, allowing for an overall view of data slices or a more detailed inspection categorized by error types ( R3). The provided color encoding choices include Slice Name, Slice Accuracy, Slice Confidence, and Spuriousness Probability, adaptable to user requirements.
The Slice Table (Fig. 2  B) presents various slice performance metrics to facilitate dataset navigation. Available metrics include Accuracy, Confidence, and Spuriousness Probability, which is produced by the Spuriousness Propagation method upon user annotation (Sec. III-E). This table assists in detecting intriguing patterns, like high accuracy with wrong attributions possibly indicative of model biases. On the other hand, low accuracy with correct attributions indicates the potential existence of mislabeled data ( R3, R4).
Our system’s third component is the Attribution Mosaic, rendering the dominant visual patterns of each slice with respect to the model’s decision ( R2, R3), either in a unified form (Fig. 2  C) or segregated by the model’s confusion matrix (Fig. 2  E). The design method for this view is detailed in Sec. III-E. Moreover, Attribution Mosaic incorporates visualization of user-specified metrics through the coloration of mosaic boundaries, providing a vital guide in pinpointing problematic slices ( R3). Annotation of a slice is facilitated by a double-click action on a mosaic tile, allowing for effortless insight collection ( R4).
The Slice Detail View (Fig. 2  D) showcases individual image samples of the selected data slice, rendering either images or attribution heatmaps. This view enables an in-depth examination of each slice ( R2).
Figure 3: Attribution-weighted feature vector generation. An image is forwarded through CNN, where the corresponding feature vector $F$ and attributions heatmap $A$ can be extracted. Through linear interpolation of $A$ to the latent space, we obtain the weight matrix $W$ to calculate the attribution-weighted feature vector $F_{W}$.
III-D Explainable Data Slice Finding
In this section, we delineate our Explainable Data Slicing method. Our approach presumes a typical CNN model architecture, comprising a sequence of convolutional layers succeeded by a fully connected (FC) layer, a commonality across most CNN models [24, 25, 12].
The process unfolds in three phases: initially, we create attribution-weighted feature vectors utilizing the interpolated attribution masks in the latent space. Subsequently, we derive an attribution representation space that maintains neighbor consistency in terms of model attributions and compute data slices. Lastly, we employ the Feature Inversion [23] to distill shared visual patterns per slice and visualize these patterns alongside model performance indicators in Attribution Mosaic.
(1) Attribution-Weighted Feature Vector Generation
In this section, we illustrate the process of generating attribution-weighted feature vectors based on attribution heatmaps ( R1, R2). It is notable that AttributionScanner accommodates interchangeable attribution explanation techniques, such as CAM [11], GradCAM [12], or RISE [14] ( R2). Our prototype leverages GradCAM due to its demonstrated efficacy via sanity checks [26] and its prevalent usage for model validation in the ML community [5, 17, 27, 28, 29, 30, 31].
As shown in Fig. 3, an input image is processed by the model to extract the feature vector $F$, and a back-propagation is conducted from the model’s predictions to obtain the GradCAM attribution masks. While a conventional approach is to depict the mask as a heatmap for the instance-level explanation, we augment its explanation capability to the latent space to foster a more comprehensive understanding of the model’s attribution patterns. Specifically, we apply linear interpolation to map image-sized masks into the latent space, yielding a weight matrix $W\in\mathbb{R}^{m\times n}$. Then the normalized $W$ with $\Sigma_{i=1,j=1}^{i=m,j=n}W_{ij}=1$ is used to compute the weighted average of $F$. The resulting attribution-weighted feature vector $F_{W}\in\mathbb{R}^{1\times 1\times d}$ can be mathematically expressed as:
$F_{W}=F\odot W=\Sigma_{i=1,j=1}^{i=m,j=n}F_{ij}W_{ij}.$ (1)
Figure 4: Comparison of the representation spaces for the hair color classification model. (a) Feature representation space computed on original feature vectors. (b) Attribution representation space computed on attribution-weighted feature vectors.
(2) Data Slice Identification
To derive meaningful data slices exhibiting similar attribution patterns ( R1, R2), we construct an attribution representation space using the attribution-weighted feature vectors. Fig. 4 provides a qualitative comparison of the feature representation space (derived from the original feature vectors) and the attribution representation space (derived from the attribution-weighted feature vectors). For both spaces, we utilize UMAP [32] for dimensionality reduction to the 2D space. At each representation space, three points are randomly sampled and their nearest neighbors are obtained to examine the consistency of their corresponding attribution heatmaps.
Figure 5: Attribution Mosaic generation. After acquiring attribution-weighted feature vectors ($F_{W}$), we compute data slices within the attribution representation space. Then Feature Inversion is conducted according to $F_{W}$ and the mosaic boundary of each slice to visualize their common patterns, forming Attribution Mosaic.
As showcased in Fig. 4, it is obvious that the neighbor instances’ model attributions from the original space are not similar, while our constructed attribution representation space has greater local consistency concerning model attributions. Therefore, our attribution representation space provides a solid basis for computing interpretable data slices with coherent attribution patterns ( R2).
On top of our constructed attribution representation space, we proceed by clustering the samples with K-Means to obtain data slices. Similarly to DOMINO [3], we apply over-clustering by incrementing the number of clusters and monitoring the model attribution consistency within each slice, until a coherent grouping of samples is attained. This measure helps prevent the omission or overlooking of rare slices, ensuring a comprehensive slice identification.
III-E Slice Summarization and Annotation
(1) Attribution Mosaic
Although we have generated slices with consistent model attributions, validating the attribution correctness of each slice necessitates individual inspection of their encompassed data samples. This process is labor-intensive and time-consuming, posing the risk of overlooking or misidentifying critical issues. Therefore, a more efficient method to encapsulate the essence of each data slice is imperative. In response, we introduce Attribution Mosaic, a novel visualization technique that elevates Feature Inversion to visually summarize data slices in a mosaic depiction ( R3). We present the pipeline in Fig. 5. First, after the execution of data slice finding (Sec. III-D), we compute the convex hull of each slice based on the projected points in the 2D space. We then distill each slice’s predominant visual patterns via Feature Inversion, integrating the shape constraint of the convex hull. The final landscape featuring slices’ attribution summaries is presented as Attribution Mosaic.
Slice Summary. A pivotal step in generating Attribution Mosaic involves visually summarizing the significant pattern in each slice ( R3). As discussed in Sec. II-B, Feature Inversion visualizes common patterns across a group of feature vectors [33, 34, 35]. We utilize this technique on the attribution-weighted feature vectors of each slice to visualize their shared attributions. Additionally, we enhance the quality of synthesized images by initializing them with the average of the original images instead of Gaussian noises [34] and parameterizing them in the Fourier space [35], making them better aligned with natural images rather than less-interpretable artificial patterns. The optimization process is defined by:
$x^{*}=\operatorname{arg\,max}_{x\in\mathbb{R}^{W\times H\times C}}\left(\ell(\phi(x),\phi_{0})+\lambda R(x)\right),$ (2)
where $\phi:\mathbb{R}^{W\times H\times C}\rightarrow\mathbb{R}^{d}$ is the representation function; $\phi_{0}$ is a target feature activation; $\ell(*)$ is the loss function where the Euclidean Loss is used here, and $R(x)$ is a Total Variation (TV) regularization term to improve the quality of results [23]. In our scenario, we designate the optimization target, $\phi_{0}$, to be the average of attribution-weighted feature vectors of a data slice, and assign the representation function to the model’s convolutional modules. We obtain an image $x^{*}$ to minimize the loss function, thereby visualizing the shared attribution patterns within a data slice from the model’s perspective.
Mosaic Boundary Generation. The main purpose of this stage is to devise a clear layout enabling effortless exploration of diverse attribution patterns across slices. Merely obtaining and displaying the Feature Inversion results over the projected cluster centroids leads to overlapped visualizations, hindering effective user exploration. We design a mosaic layout, optimized for space-efficiency in information representation. Specifically, we compute convex hull [36] of each slice in the attribution representation space, yielding a convex hull for every slice with no overlap amongst them.
Mosaic drawing. Given that the convolutional modules take rectangular-shaped input, the original Feature Inversion algorithm solely generates rectangular visualizations. To preserve equivalent visualization in a mosaic form, we use our generated mosaic boundary as a constraint for optimization. Namely, we define the shape of $x^{*}$ as the minimal rectangle encompassing the corresponding convex hull, and in each iteration, we zero out pixels outside the hull, compelling the optimization to concentrate on the interior pixels. After rendering all feature visualizations, they are exhibited at their respective positions within the Attribution Mosaic.
(2) Slice Annotation and Spuriousness Propagation
The Attribution Mosaic facilitates rapid comprehension of the main content of different slices, supporting the identification of potential errors in data or models ( R4). A prevalent and harmful issue in ML models, highly visible through Attribution Mosaic, is “spurious correlation”, where a model erroneously utilizes irrelevant image features for classification. For instance, in hair color classification, a model should make decisions according to hair features, termed as “correct/core correlation”, while the use of other features (e.g., background or facial attributes) denotes “spurious correlation.”
Such spurious correlations, regardless of model accuracy, pose substantial challenges including poor generalization in production [37] and AI fairness issues [38]. We introduce a metric termed Spuriousness probabilities to streamline the detection of spuriousness, ranging from 00 to $1$, indicative of the likelihood of spurious correlation existence. We employ the Label Spreading method [39] to propagate user annotated spurious/core information to other slices. Specifically, the minimal input requirement from users is one slice annotation of “core” (Spuriousness=00) or “spurious” (Spuriousness=$1$), and then AttributionScanner will propagate this information to create a Spuriousness Probability for the rest slices.
Table I: System Design Considerations.
Design Components System Capabilities
Slice Table Slice Detail View Scatter Plot Attribution Mosaic Slice Performance Slice Attribution Instance Performance Instance Attribution
The Spuriousness probabilities obtained through the Attribution Mosaic offer two benefits. First, they streamline slice exploration by highlighting potential spurious correlations, supporting identifying and assessing problematic slices. Second, after users’ verification, these probabilities inform the model-side strategies for issue mitigation (Sec. III-G) ( R4).
Figure 6: AttributionScanner applied to the validation of a bird category classification model trained on the Waterbirds dataset. The user finds the correct model behavior (bird patterns) corresponding to a well-performed Slice 34 and annotates it as “core: birds” B and C. By switching the Attribution Mosaic to the confusion matrix view D and investigating the underperformed slices with accuracy sorting A, the user identifies a problematic Slice 32 that has high false negatives D1, which turns out to use spurious feature of land backgrounds (BG) to predict landbirds.
Figure 7: Example of spurious correlation mitigation with CoRM for hair color classification: 1 The Attribution Mosaic indicates a problem with a spurious feature: mouth. 2 An example of how noise is occupied by CoRM: Original Image (left), GradCAM activations (middle), noise added to the spurious regions of the image (right).
III-F System Design Considerations
In designing AttributionScanner, various configurations of visual components’ combinations have been evaluated to ensure optimal system capabilities to fulfill our design requirements discussed in Sec. III-A, as outlined in Table I. Our system should be capable of providing four-faced information: (1) slice performance indicator ( R3), (2) slice attribution overview ( R2, R3), (3) instance performance indicator ( R2), and (4) instance attribution inspection ( R2, R4). The notions in “design components” are consistent with Fig. 2.
In Table I, the initial configuration, “Slice Table & Slice Detail View,” is capable of presenting slice performance indicators and instance attributions. However, the numerical matrices of performance indicators are not sufficient in guiding users to uncover hidden model issues, such as spurious correlations that can happen regardless of high or low performance. A subsequent iteration included a scatter plot showing the attribution representation space. This configuration enables the system’s instance performance indicator capability but still misses the slice attribution overview. To the best of our knowledge, there exist few approaches that visually elucidate potential model spuriousness in scale, where our third configuration fulfills this gap — “Slice Table & Slice Detail View & Attribution Mosaic, ” which comprehensively satisfies all the essential system capabilities and fulfills our design requirements.
Attribution Mosaic, alongside the Slice Table and Slice Detail View, enriched the system’s ability to provide granular insight into both slice and instance-level attributions and performance indicators ( R2, R3). The coordinated Slice Table and Attribution Mosaic can guide users to pinpoint hidden issues, and the Slice Detail View enables users to further verify their insights ( R4). This robust configuration underscores our meticulous design process in assuring that AttributionScanner is not only capable of furnishing critical insights but also does so in an intuitive and user-friendly manner.
The tabulated design considerations and the ensuing choice of system component configuration aim at maximizing the system’s utility and user-centric functionality. With our finalized configuration (the last row of Table I), users are empowered with sufficient support for problem detection, interpretation, and mitigation in the model validation.
III-G Slice Error Mitigation
One effective approach to mitigating spurious correlations in ML models is through model re-training, which can improve the model’s robustness without changing its architecture. We leverage the Core Risk Minimization (CoRM) method introduced in [37]. CoRM corrupts non-core image regions with random Gaussian noise and retrains the model using the noise-corrupted data, which has been shown to be effective in mitigating a model’s reliance on spurious features.
[…]
Figure 8: Usage patterns for AttributionScanner derived from case studies. The analysis can start with A a rank-driven exploration (ranking the slices by A1 score (e.g., model accuracy) or A2 spuriousness), or B a visual-driven exploration. After inspecting slice summary in Attribution Mosaic ( B), they can inspect individual samples in the Slice Detail View ( C). Finally, they can annotate their insights ( D).
IV Case Studies
In this section, we present two case studies with publicly available vision datasets to benchmark and evaluate the capabilities of AttributionScanner. The primary objective of these case studies is to demonstrate how AttributionScanner empowers ML experts and practitioners to detect, evaluate, and interpret potential issues in vision models. Fig. 8 illustrates common usage patterns of our system: users can begin the analysis with either a rank-driven evaluation A or a visual-driven evaluation  B. Once a slice of interest is identified, users can delve deeper into individual samples in the Slice Detail View  C. Finally, the user can annotate the data slices and continue their analysis  D.
IV-A Hair Color Classifier Validation - Finding Edge Cases
$\bullet$ Overview. This case study involves the Large-scale CelebFaces Attributes (CelebA) dataset [40] with $202,599$ number of face images. The label of each image is one of {not gray hair, gray hair}, refer to labels $\{0,1\}$, respectively. With the train, validation, and test splits of $(8:1:1)$, we adopt transfer learning to train a ResNet50 [25] binary image classifier. After iteratively fine-tuning hyper-parameters, we obtain a model with $98.03\%$ classification accuracy. The ML experts would then explain and troubleshoot this hair color classifier with AttributionScanner based on these settings: n_neighbors $=5$, min_dist $=0.01$, and n_components $=2$ for the UMAP algorithm, and n_clusters $=50$ for K-Means.
Figure 9: Examples of findings when ML experts validate a hair color classifier, where experts discover slices corresponding to the model’s core/spurious correlations and wrong labels.
$\bullet$ Does the model behave correctly on well-performing slices? One basic expected behavior for a hair color classification model is to catch hair features. At first glance of the Attribution Mosaic (Fig. 2  C), ML experts notice slice_22 and slice_5 on the left, which lie separately with others on the Attribution Mosaic. Their mosaic visualizations indicate gray hair patterns, which means the model uses the correct features, i.e., core features. By verifying the corresponding attribution heatmaps (Fig. 2  D), they confirm the correctness of this insight and annotate both slices as core feature with description Core: gray hair. Upon experts saving the annotation, AttributionScanner automatically propagates the annotation and provides a Spuriousness probability of each slice. With this guidance, the experts observe many slices lying on the right of Attribution Mosaic have higher Spuriousness probabilities (Fig. 2  C  E). Their mosaic visualizations do not show any hair patterns, leading to valid doubts that the model does not behave correctly on such slices. Besides, they also notice that the model has correct predictions on these slices (with $100\%$ prediction accuracy). This makes them worry that the model is largely biased by spurious features. Through investigation, they find the model mistakenly utilizes mouth and eyes to predict hair color for those top-performed slices, such as slice_25, slice_14, and slice_2 (Fig. 9).
$\bullet$ Why slices underperform? The experts are also eager to investigate the underperformed slices. By sorting the Slice Table (Fig. 2  B) by ascending order of accuracy, they notice the model only achieves $72.41\%$ accuracy on slice_44 and find that the slice’s visualization only shows colorful patterns without any recognizable content. After checking the attribution heatmaps, they find that the model mistakenly uses image backgrounds to make hair color predictions (Fig. 9). This spurious correlation issue stands out, and they annotate this slice as “spurious” with the description “Spurious: backgrounds”. Similar issues also occur in its neighborhood slices via Attribution Mosaic, and AttributionScanner automatically propagates them with higher Spuriousness.
Figure 10: Examples of findings when ML experts validate a bird category classifier, where experts identify problematic slices where the model uses backgrounds (BG) to classify birds.
$\bullet$ What underlying factors contribute to unexpected behaviors? ML experts are interested in understanding why such unexpected model behaviors happen. By sorting the Slice Table by descending order of Spuriousness, they get a list of slices with high Spuriousness possibilities. They switch Attribution Mosaic into the confusion matrix form (Fig. 2  E) to study further details. By clicking on the name of “slice_35”, they highlight this slice on the four sub-views and inspect the provided explanations, where they notice that images that lie in the “FN” group of this slice have wrong labels — they should be labeled as “not gray hair” rather than “gray hair”. They mark this issue as wrong labels and wonder whether this situation is common. By investigating its neighborhood slices on Attribution Mosaic, they find and mark another slice with wrong labels through a few clicks, slice_2_FN.
$\bullet$ Insights summary. Fig. 9 presents examples of the detected issues with the current model, which include wrong attributions in underperforming slices, model biases in well-performing slices, and noisy label issues. Through the aforementioned procedure in this case study, we demonstrate how AttributionScanner supports users to uncover and interpret potential model issues with visual summaries and other guidance. Based on annotations from ML experts, we employ the CoRM framework to mitigate the detected errors, which will be evaluated in Sec. V.
IV-B Bird Category Classifier Validation - Detecting Bias
[…]
While in this study, ML experts are aware that the model is likely biased by backgrounds — using water (land) background to classify waterbirds (landbirds). However, such priori knowledge is hard to establish in real-world applications because of the scarcity of additional well-labeled metadata. And hence the experts assume such information is unknown and want to validate whether AttributionScanner can make the potential model biases stand out by only utilizing the original images and the trained model.
$\bullet$ Does the model exhibit bias? To answer this key question, experts start by investigating the underperformed slices. From the Slice Table (Fig. 6  A), they sort slices by ascending order of accuracy and select the worst-performed slice_38. The coordinated information provided by Attribution Mosaic and model attribution heatmaps highlights a spurious correlation problem—the model uses water backgrounds to classify birds (Fig. 10). The experts annotate this slice as “spurious”, and AttributionScanner automatically propagates this annotation. They verify the propagation correctness on neighborhood slices and annotate slice_41 as “spurious” with “water backgrounds” (Fig. 10). Moreover, the experts identify an underperformed slice_32 that is not clustered with the current ones and is given a high Spuriousness possibility. They investigate it and verify that the model utilizes “Spurious feature: land backgrounds”, to predict bird classes. Through the Attribution Mosaic in the confusion matrix form (Fig. 6  D), they find such spurious correlations result in many false negatives (Fig. 6  D1), where the model uses land backgrounds to mistakenly predict many “waterbirds” as “landbirds”.
Table II: Quantitative evaluation of the performance of the hair color classification models trained on the original and AttributionScanner(AS)-improved CelebA dataset.
Training Procedure Clean Acc ($\uparrow$) Core Acc ($\uparrow$) Spurious Acc ($\downarrow$) RCS ($\uparrow$)
Baseline 98.02688 97.21170 97.61917 0.067720
AS (Annotation) 98.02688 98.02185 96.92958 0.216351
AS (Propagation) 98.08225 98.16278 96.01852 0.368512
$\bullet$ Is the detected bias prevalent across all slices? Why or why not? ML experts are interested in determining whether the detected bias is pervasive throughout the dataset. By analyzing slices that are distant from the annotated ones in the Attribution Mosaic and are assigned with low Spuriousness possibilities, they discovered that the farthest neighbors, namely slice_34 and slice_5, correspond to core features. This suggests that the model can correctly capture bird regions in these slices (Fig. 6  B  C), which raises a follow-up “why” question. To understand in what circumstances when the model fails. They browse the original images from slice_32 (spurious feature) and slice_34 (core feature), respectively, as shown in Fig. 6  D and  C. They find slice_32 has very similar land backgrounds and very different birds, while on the other hand, the birds’ appearance of slice_34 is very consistent. This finding explains why this biased model can successfully capture the core features from slice_34 but fails at slice_32—greater similarities in the representation space indicate stronger features. In other words, core features in slice_42 are strong enough to support the model’s robustness against bias. Such insights are helpful in improving model robustness and have been further studied by ML experts [43].
$\bullet$ Insights summary. A summary of insights obtained from this case study is provided in Fig. 10, where ML experts validate the existence of model biases and extract slices corresponding to different biases. In the following Sec. V, we evaluate whether AttributionScanner can mitigate model errors by incorporating human feedback.
Figure 11: Qualitative evaluation with GradCAM attribution heatmaps. We visually compare the original and the AttributionScanner-improved models for hair color and bird category classification, respectively. In each sub-figure, the first row refers to the original model (baseline), and the second row refers to the improved model (ours). We can observe AttributionScanner suppresses the spurious correlations by using the correct features for predictions.
Table III: Quantitative evaluation of the performance of the bird category classification models trained on the original and AttributionScanner(AS)-improved Waterbirds dataset.
Training Procedure Clean Acc ($\uparrow$) Core Acc ($\uparrow$) Spurious Acc ($\downarrow$) RCS ($\uparrow$)
Baseline 85.73812 82.98582 82.82901 0.004878
AS (Annotation)
89.57465 86.40534 79.81651 0.195062
AS (Propagation)
90.40867 87.40617 77.89825 0.274038
V Evaluation
In this section, we conduct both quantitative and qualitative evaluations to validate whether AttributionScanner can indeed leverage human insights to enhance vision models’ performance while reducing their reliance on spurious features.
V-A Quantitative Evaluation
The quantitative evaluation involves four matrices introduced in [37], including clean accuracy, core accuracy, spurious accuracy, and relative core sensitivity (RCS). We revisit their definitions as follows:
$\bullet$ Clean Accuracy. The original model accuracy, where larger values are indicative of better overall accuracy.
$\bullet$ Core Accuracy $acc^{(C)}$. The model accuracy when spurious regions are occupied with Gaussian noise, where larger values are indicative of the model’s more reliance on core regions.
$\bullet$ Spurious Accuracy $acc^{(S)}$. The model accuracy when core regions are added with Gaussian noise, where larger values are indicative of the model’s more reliance on spurious regions.
$\bullet$ RCS. This matrix quantifies the model’s reliance on core features while controlling for general noise robustness. It is defined as the ratio of the absolute gap between core and spurious accuracy, and the total possible gap for any model between core and spurious accuracy. Represented as $RCS=\frac{acc^{(C)}-acc^{(S)}}{2\times min(\alpha,1-\alpha)}$, where $\alpha=\frac{acc^{(C)}+acc^{(S)}}{2}$. RCS ranges from 00 to $1$, a higher value indicative of better model performance.
In the two case studies discussed in Sec. IV, ML experts annotated five spurious slices in the CelebA dataset, {slice_25, slice_14, slice_2, slice_44, slice_26}, and six spurious slices in the Waterbirds dataset, {slice_38, slice_41, slice_32, slice_29, slice_4, slice_10}, respectively. For each case study, AttributionScanner automatically ran the Label Propagation algorithm and exported both the users’ annotation records and the propagated Spuriousness for further investigation.
To thoroughly evaluate our introduced method, we validate three models for each case. The models marked as “baseline” are the original trained models obtained at the beginning of each case study. The models marked as “AttributionScanner” are re-trained using the CoRM method after adding noise to “spurious” slices according to results exported from AttributionScanner. In particular, “Annotation” indicates that only user-annotated spurious slices were corrupted with noise, while “Propagation” indicates that propagated Spuriousness is used to identify spurious slices to be added with noise.
Table II and Table III present the quantitative evaluation results for the two case studies, respectively. Our presented AttributionScanner can significantly improve the vision model’s overall performance with reduced spurious correlations. Furthermore, our Label Propagation largely reduces human effort by automating the annotation process, and achieves the best performance in this quantified evaluation. Overall, our results demonstrate that AttributionScanner is an effective approach for mitigating spurious correlations in machine learning models, and the Label Propagation algorithm is a valuable tool for automating the annotation process.
V-B Qualitative Evaluation
Results in Sec. V-A highlight the best performance of AttributionScanner equipped with Label Propagation in mitigating spurious correlations. For further evaluation, we visually compare models’ attributions via GradCAM in Fig. 11. Refer to Sec. IV-A, the hair color classifier originally has spurious correlations dominant in {slice_44, slice_25, and slice_14}, where the model uses image backgrounds, mouth, or eyes to predict hair color, respectively. With the help of AttributionScanner, the model’s misattribution is successfully fixed, directing focus towards the correct hair regions (refer to Fig. 11(a)). As for the bird category classification model, it originally focuses on spurious features, water/land backgrounds, to decide whether there are water/land birds on the input image {slice_38, slice_41, and slice_32} (refer to Sec. IV-A). In Fig. 11(b), we can observe that AttributionScanner mitigates these issues by helping the model to focus on the core bird features.
V-C Experts Feedback
We conduct two case studies (see Sec. IV-A and Sec. IV-B) with ten ML experts to evaluate AttributionScanner. None of them are co-authors of this paper, and they had not previously seen AttributionScanner. Five are male and five are female. Six of the experts have over five years of experience in ML, two have between three to five years of experience, and two have between one to three years of experience. All experts need to conduct model validation in their research and have experience working with vision models. Our findings are drawn from their comments during the study, as they follow the “think-aloud” protocol, and additional feedback is gathered through exit discussion and questionnaire.
Summary of Likert-Type Questions.    The exit questionnaire included nine Likert-type questions on a seven-point scale, aiming to evaluate AttributionScanner from various aspects. The feedback, as shown in Fig. 12, reflects a positive overall impression. Notably, in the usability evaluation, five experts strongly agreed and three agreed that our system is user-friendly and understandable. In terms of effectiveness, eight experts strongly agreed that our system aids in validating models. All experts expressed strong confidence in the insights provided. Considering that Feature Inversion (FI) results are synthesized images that might initially be hard to grasp, we included questions to assess this design aspect. Seven experts strongly agreed, and two agreed that FI is helpful for understanding slice patterns. While two experts expressed neutral or somewhat disagreeing opinions about understanding slice contents based solely on FI, eight strongly agreed that AttributionScanner’s coordination of FI and other views supports a comprehensive understanding of slice contents. Lastly, all experts expressed interest in using our system in the future and would recommend it as a valuable tool for model validation.
Interpretable Model Validation.    All experts acknowledged that AttributionScannereffectively helps ML model validation”. They remarked that the complimentary visual summaries and attribution heatmaps help them “easily understand slice issues”. “It’s often hard to figure out what’s going wrong with a model”, one expert said, “This system guides me to find where I should look and explains the issues intuitively.” Another expert noted, “The ability to browse what’s happening in different slices at once greatly helped me understand the model.” They commented that they are confident about their insights because our system “highlights slice errors and provides supportive evidence”. Three of them are eager to see what issues AttributionScanner could uncover for the models they are currently using.
Attribution Mosaic.    All participants showed particular interest in the Attribution Mosaic and asked about how it was designed in the interview. Seven experts were impressed by its ability to summarize model attributions across data subgroups. Two experts noted that “it highlights model issues intuitively.” Another described it as “impressive and novel,” emphasizing that it “highlights and explains the model’s failures.” Another expert commented, “This view provides a new angle for me to identify and understand the data/model issues.” All experts quickly adapted to using Spuriousness propagation and utilized the Spuriousness matrix to speed up their annotation. Specifically, five experts found that the Spuriousness propagation “surprisingly helpful,” and six remarked that it “really saved my effort”.
Other comments.    Our experts provided feedback on additional aspects. One expert suggested that we directly display predictions and labels via text in each confusion matrix panel. “It would be better if you show ‘label: waterbirds; prediction: landbirds’ rather than ‘TN’”. Another expert discussed the potential of applying our system to validate multi-classification models. Following our brief explanation of how AttributionScanner produces results based on the model’s predicted class without restrictions on the number of classes, the expert commented, “Your system seems to be fully adaptable. And I would love to see this implementation in future work.”
Figure 12: Expert perception of the system, according to eight Likert-Type Questions using a 7-point scale. (AS = AttributionScanner.)
VI Discussion and Future Work
Potential risks introduced by XAI techniques.     In this work, we leverage state-of-the-art XAI techniques including GradCAM and Feature Inversion, where the first is one of the most widely-adopted attribution-based explanations for neural networks [5, 44, 45, 46] and the second is an advanced technique to visualize what a model is looking for [33, 47, 48]. However, there has been continuing discussion on the reliability and trustworthiness of XAI techniques [26, 49], indicating the potential risks of using them. Even though both of them have been proven to retain a level of correctness through validations such as saliency checks [26], we are aware of such risks and involve human auditions in our workflow to alleviate them. In the future, we plan to design more trustworthy model explanations by reasoning about the causality relationships in model decisions [50, 51], and continue to involve human-in-the-loop to mitigate pitfalls in AI-automated tools.
Propagation approaches for the hypothetical spuriousness.     We adopt the Label Propagation method [39] to automatically spread the slices’ spurious annotations from users to other slices as discussed in Sec. III-E. This is under the assumption that the distance between the slices corresponding to spurious correlations (e.g., water backgrounds) and the slices corresponding to core correlations (e.g., birds) is large. However, such assumptions can be invalid in other scenarios. For example, an image classification model trained on ImageNet [52] includes various different classes such as “husky dog” and “king penguin”. It is likely that two slices far away from each other are both corresponding to the core correlations. To address this, we plan to incorporate class similarities to the Label Propagation matrix to improve the precision of the generated hypothetical Spuriousness.
Better guidance in data slice finding.     We include the model’s performance indicators such as the classification accuracy and the average prediction confidence, as well as the Spuriousness probabilities in the Slice Table of our system to guide users to find interesting slices (refer to Fig. 2  B). Although the current guidance is proven to be effective in assisting users to detect slice issues, we plan to make further improvements. We aim to introduce new metrics calculated based on the model’s performance, the slice pattern similarity, and the Spuriousness. By doing this, we will provide more informed guidance to our users by leveraging human and model’s knowledge simultaneously, further improving the effectiveness of our model validation workflow.
VII Conclusion
In this work, we present AttributionScanner, a novel human-in-the-loop model validation system empowered by metadata-free data slice finding and XAI techniques, which can support effective detection, explanation, and mitigation of potential errors and biases in vision models. To achieve this, we first present a metadata-free data slice finding method by using attribution-weighted feature vectors to group unstructured image data into data slices with consistent model attributions. Also, we design a slice pattern summarization view, Attribution Mosaic, which visually presents the shared significant patterns of each slice attributing to the model’s decision. This way, AttributionScanner is able to provide explainable data slices that reveal critical model problems, such as spurious correlations and mislabeled data. We also demonstrate the effectiveness of this work in performance improvement and bias mitigation through both quantitative and qualitative evaluations. To sum up, without additional meta-information, multi-modal model, or other background information, AttributionScanner enables an effective model validation and error mitigation workflow to alleviate underlying issues in ML models. We hope this work can inspire future research in human-AI teaming to improve AI trustworthiness and accountability.